When 250.0 mg of eugenol was added to 100.0g of camphor, it lowered the freezing point of camphor by 0.620C. Calculate the molar mass of eugenol.
1
Expert's answer
2017-11-30T15:12:06-0500
ΔT = Cm · Kf
Сm = n/m n = m/M M (C10H12O2) = m/n
Kf (camfora) = 4.0
Cm (C10H12O2) = ΔT/Kf = 0.62/4 = 0.155 m
n (C10H12O2) = 0.155 · 100 = 15.5 mol
M (C10H12O2) = 250 / 15.5 = 161 g/mol